jan 15,2000 evening shift


Subject: jan 15,2000 evening shift
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:26:13 CST


Shift Summary for eve shift, 01/15/00

On Shift: Mike A., Ricardo B.

Hours of beam delivered at all intensities: 8.5
Hours of beam delivered at nominal intensity:8.0
Hours of E799 data (special runs not included): 7.8
Hours of special runs (laser scans, muon runs...):none

Accounting of lost time: fresh start after ktev2 stopped writing,
 pipeline resets, kfrends

===============================================================

Runs ended this shift:

Run 15540 - 11M events - Beam_799

===============================================================

Tapes brought to FCC this shift: none

Accesses during shift:none

===============================================================

Description of Shift: taking regular E799 data

run spill time comment
----------------------------------------------------
15540 15:40 spill structure bad, complain to MCR, they
                     reduce intensity to 4E12 while working on it.
                     reset of QXR fixes it

             16:15 spill OK now, intensity back to 9E12.

             16:32 pipeline errors (ordinal mismatch) KFREND

             17:20 ask for 9.5E12, they are actually almost there
                     already.
          
             21:10 "check L3 accept fraction", Notice that no
                     events are being written on ktev2 drives,
                     although triggers are running and memories
                     are working normally. Call Ed, he can't tell
                     what's wrong either. Stop run, it stops normally.

15541 21:30 new beam_799 run after DRT_FRESH_START, no problems,
                     data appends to tapes normally.

             21:10 TRD alarm, TRD2 back plane drew higher than
                     normal current for about 10 min then back to
                     normal.

15541 ~60-80 23:00 pipeline errors, reset several times and the
                     problem comes back, finally call Bob T. He
                     suggests reset with DBC, After about 23:15 no
                     more errors.
                23:50 Bob T calls back and says his analysis of event
                     dumps shows that all the problems were in fact
                     not pipeline errors but rather ordinal mismatches
                     in stream 5.
----------------
 
 For next shift:
         Ed wanted to be paged at next run change to install new
  CsI constants.

         We are getting off and on Epicure errors from slow DAQ,
 NM2S1 is reading back slightly lower than tolerance.

      



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:26:15 CST